Give an example of a regular space that is not normal
1
Expert's answer
2020-05-13T18:01:47-0400

An example of a regular space that is not normal is the Sorgenfrey plane - "\mathbb{R}_l\times\mathbb{R}_l"   

 It is regular because it is the Cartesian product of regular spaces. It is not normal because any subset 

"A" of "-\Delta=\{x\times(-x)|x\in\mathbb{R}\}"  is a closed subspace of "\mathbb{R}_l^2" and it can be shown that there do not exist disjoint open sets about "A" and "-\Delta \setminus A" in "\mathbb{R}_l^2" .